OpenStudy (anonymous):

blank sqrt over 3 = 4? explain.

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

well now it equals 2 lol

OpenStudy (gorv):

\[\sqrt{3}=4\]

OpenStudy (gorv):

???

OpenStudy (anonymous):

blank sqrt over 3 = 2 lol

OpenStudy (gorv):

divide both side by sqrt 3

OpenStudy (anonymous):

so i do sqrt of 3 then divide that number into 2 and 3 seperately?

OpenStudy (anonymous):

im lost.....

OpenStudy (gorv):

divide both side

OpenStudy (anonymous):

is it like what i said a few minutes ago??

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

sqrt of 3 is 2 sooooo 2 divided by 3 that doesnt make sense

OpenStudy (anonymous):

2 divided by 2 is 1 thats all i know so 1/3?

OpenStudy (gorv):

\[blank*\sqrt{3}=2\]

OpenStudy (gorv):

\[\frac{ blank*\sqrt{3} }{ \sqrt3 }=\frac{ 2 }{ \sqrt3 }\]

OpenStudy (anonymous):

ohhh okay that makes sense now

OpenStudy (anonymous):

but its suppose to equal 2 lol

OpenStudy (anonymous):

i cant put a fraction

OpenStudy (gorv):

we divided on both side left and right

OpenStudy (gorv):

so what u will get??

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

oh 2/3 = 2?

OpenStudy (gorv):

how u got it ??

OpenStudy (gorv):

on left sqrt 3 will get cancelled

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

waiy

OpenStudy (gorv):

u got it now ??

OpenStudy (anonymous):

hmm not really

OpenStudy (gorv):

what u missed??

OpenStudy (anonymous):

cause you cant really divide 3 into anything but 6 really

OpenStudy (anonymous):

so is it 2?

OpenStudy (anonymous):

lol i cant get this question wrong ive already missed 2....

OpenStudy (anonymous):

@aliboo997

OpenStudy (gorv):

we are not dividing 3 into anything

OpenStudy (gorv):

\[blank*\sqrt3=2\]

OpenStudy (gorv):

it is your Q

OpenStudy (gorv):

so we need to make the word blank alone to find its value

OpenStudy (gorv):

but sqrt3 is multiplied to it so we need to remove it

OpenStudy (gorv):

so to remove it we divide both side by \[\sqrt{3}\]

OpenStudy (gorv):

which cancel or cut or remove sqrt on left side and make blank alone

OpenStudy (anonymous):

well i closed the test i have to get 8 out of 10 and i missed 3 when i can only miss 2 idk

OpenStudy (gorv):



OpenStudy (anonymous):

im still lost but oh well.... i have a new problem :d

OpenStudy (anonymous):

-17 sqrt 11 - 5 sqrt 11 i got -22 sqrt 3 and it said that was wrong how??